a ferry leaves a harbor and travels to a resot island at an average speed of 18 mph. On the return trip the ferry travels an an average speed of 12 mph because of fog. The total time for the trip is 6 hours. How far is the isalnd from the harbor.
Answer by ewatrrr(24785)


Hi
D = r*t OR D/r = t
Questions states*** Total time was 6 hrs

D/18mph + D/12mph = 6hrs |multiplying each term by 36mph so as all denominators=1
2D + 3D = 6hr*36mph
5D = 216mi
D = 43.2mi,the distance of the island from the harbor